OpenStudy (anonymous):

If the point (1, 3) lies on Quadratic 1, and (1, 4) lies on Quadratic 2, which of the following points must lie on the polynomial that is the sum of these two quadratics?

OpenStudy (anonymous):

I need the choices :)

OpenStudy (anonymous):

the question is asking which point will lie on the same line and be a sum of the two. since both points have x being 1. you know that your x value must be 1 or else it wont be on the line. 3+4 is 7. which means that the point is (1,7)

OpenStudy (anonymous):

Parth, your 12. Stop showing off lol

OpenStudy (anonymous):

hmmm

Parth (parthkohli):

Hmm I was wrong. I don't know what the question is.

OpenStudy (anonymous):

look at @annas answer, it is correct

OpenStudy (anonymous):

Parth tried his best instead of copy pasting from yahoo answers. I appreciate him

Parth (parthkohli):

What is meant by 'Quadratic 1' and 'Quadratic 2'?

OpenStudy (anonymous):

actually all the answers are a bit odd \(p_1(1)=3,p_2(1)=4, p_1(1)+p_2(1)=3+4=7\)

Parth (parthkohli):

 Can anyone explain what the question said?